OpenStudy (yacoub1993):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept

OpenStudy (yacoub1993):

@Directrix

Directrix (directrix):

The y-intercept is the point where the graph crosses the y-axis. Every point on the y-axis has x-coordinate zero. So, to find the y-intercept of y = -2x - 3, substitute zero in for x and solve for y. y = -2x - 3 y = -2*0 - 3 y = ?

Directrix (directrix):

@Yacoub1993

OpenStudy (yacoub1993):

-3 but i still didnt understand

Directrix (directrix):

Here is your line. You just find the coordinates of the point where the line y = -2x -3 cuts the y-axis of the plane.